OpenStudy (anonymous):

Hi. I am having some trouble making sense out of something from my book..1<=2cosu<=2. When they solve for theta they say the bounds are 0<=u<=pi/3 which seems wrong to me.

OpenStudy (anonymous):

Sorry that was suposed to be a 2

OpenStudy (phi):

are you sure it is \[ 1 \le 2 \cos u \le 2 \] they start with ?

OpenStudy (anonymous):

Yeah I sure think so. I am parameterizing a sphere x^2 +y^2 +z^2=4 with the constraint 1<=z<=2. When I paramaterize sphere I get r(u,v)=<2sinucosu,2sinusinv,2cosu>. Since 1<=z<=2 and Z=2cosu. 1<=2cosu<=2

OpenStudy (phi):

if so, divide by 2 to get \[ \frac{1}{2} \le \cos u \le 1 \] take the inverse cos of both limits to get pi/3 and 0 which means u >=0 and u<= pi/3

OpenStudy (anonymous):

So isn't that pi/3<=u<=2pi?

OpenStudy (anonymous):

just don't understand how it seems the direction changes?

OpenStudy (anonymous):

Do the < change direction when taking inverse cos?

OpenStudy (phi):